The Wanderers Guide to the World Wide Web^{∗}

Revision 0.1 - 19 January 1994

1 What is the WWW?

The World Wide Web (WWW, or W3) is a wide-area hypermedia information retrieval initiative aiming to give access to a large universe of documents from all over the Internet. The project was originated at CERN and collaborated upon by a large, informal, and international design and development team. The WWW provides a standard, easy way to obtain information by unifying HTTP, WAIS, Gopher, FTP, telnet, and other protocols. WWW servers in many countries allow users to browse hundreds of thousands of documents on every topic, display maps and pictures, play sound files^{1}, watch animated movies, and more. Hypertext documents may contain links which when selected will take you to other documents containing more information. WWW client programs provide a simple interface to access these documents without requiring the user to be familiar with the underlying protocols.

2 WWW Clients

The most popular program for accessing the Web is Mosaic. Simply do:

     athena% add sipb; Mosaic &

After a minute or two, a window will appear displaying the SIPB home page, which is an initial starting point for exploring the Web.

Another (unsupported, but still good) client is lynx, a tty-based interface that can be run over dialup, and which is started by doing:

     athena% add outland; lynx

Other clients are available for emacs, Windows, and Macintoshes. For more information on these, drop by the SIPB office, W20-557, or call x3-7788. The rest of this document will explain how to wander the Web using Mosaic.

3 Following links

Hypermedia consists of documents with links to other documents, which may contain text, pictures, sound, and/or more links. Reading becomes nonlinear: you might be reading one document, jump via a link to another document to read more in-depth information about a certain topic, and then jump back to your original document to finish.

By default, links in Mosaic are blue and underlined. Clicking with your left mouse button on a link will cause the link to be followed and the selected document to be shown. Clicking with your middle mouse button will cause the link to be followed and the document to be displayed in a new window. Several windows may be open at once for viewing multiple documents; the “Close Window” button at the bottom of a window provides a way to exit those windows you no longer need. If you visit a document and then return, the link to it will be shown in a darker blue and with a dashed underline so you can easily tell which documents you have already viewed.

Mosaic keeps a list in memory of what documents you have visited. Pressing the “Back” button will take you back to your previous document, and pressing the “Home” button will take you back to your original home page—especially useful if you get lost.

4 The SIPB Home Page

Easily recognized by the large fuzzball at the top, the SIPB home page is the default starting point for MIT users. From here, you can visit many of the more interesting places on the Web. Local weather, status of public clusters, Discuss, the MITSFS library index, and other information is readily available. Links exist to newspapers like the Freshman Fishwrap and The Tech, and to such local MIT groups as the Laboratory for Nuclear Science and the LCS/AI Reading Room. Also available are links to the home pages of various MIT students; you can even learn to make your own personalized home page and have it added to this list.

∗Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 Student Information Processing Board of the Massachusetts Institute of Technology

1On Athena, these can only be played on Sun workstations, and please use headphones so as not to disturb others in public clusters.

5 More Interesting Places

If you are just wandering around looking for interesting documents, the SIPB home page contains a link to a list of really cool things to check out.

The SIPB home page also links to the NCSA Mosaic home page and to the Mosaic demo document. These two documents provide examples further explaining how to navigate via links, describe the Web in more detail, and contain links to good starting points for exploring the vast amount of information available online.

Another way to start your search for information is to select the option “Internet Starting Points” from the “Navigate” menu at the top of the Mosaic window. This link will take you to a page containing links to documents such as other universities’ home pages. Also available are links to the Library of Congress Vatican Exhibit, White House papers, physics papers from Los Alamos National Laboratory, Internet RFCs, and an online magazine.

6 Searching for Specific Information

With hundreds of sites each providing thousands of documents, finding information that you are interested in can be challenging. To help make things easier, the Internet Starting Points document (see previous paragraph) also has listings of information organized by topic and subtopic, by various services, and by sites. A list of new places and services on the Web is also in this document.

Some documents have special search options; for example, the MITSFS library index allows you to perform an author or title search on books in their collection.

You can also search the current document for a particular word or string by selecting “Find In Current” under the “File” menu. This will pop up a dialogue box in which to specify what to search for.

7 The Hotlist

What if you find a really interesting document that you would like to visit often again, but without having to try to navigate far? You can add this document to your hotlist, which is simply a personal list of neat places. Selecting “Hotlist” from the “Navigate” menu will show your current hotlist and allow you to add the current document to it. You can then easily jump to any document in your hotlist.

8 More Useful Commands

The “File” menu at the top of the Mosaic window contains a few other useful commands. You can print the current document, save it permanently in your home directory, or mail it to a friend. The “Options” menu lets you set such things as font type and size.^{2} The “Help” menu at the top right provides assistance on a variety of topics; select “Manual” for in-depth help.

9 Problems?

The most common problem you will encounter is a document being unreachable. This could be due to a server being down, network delays, or restrictions on who may view the document. Pressing “Back” will take you to where you came from. Clicking on the spinning globe icon will stop Mosaic from downloading information; this is useful if you want to give up on trying to get a large document when the network is being slow. Finally, if you have other questions or problems that are not answered in the “Help” menu, contact someone at SIPB, W20-557, x3-7788, who will be glad to assist you. Have fun, and don’t panic!

2These features and more can also be customized in the file ∼/.Xresources.